OpenStudy (anonymous):

The shape of a roller coaster is modeled by a polynomial function, R(x). Describe how to find the x-intercepts of R(x) and how to construct a rough graph of R(x) so that the engineer can predict when there will be no change in the direction of the coaster. You may create a sample polynomial to be used in your explanations.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

got disconnected

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

For any \(\large\color{black}{ R(x)=..............}\) , to find the x-intercepts. set: \(\large\color{black}{ 0=..............}\)

OpenStudy (solomonzelman):

see what I am doing?

OpenStudy (solomonzelman):

and then you solve for x, for whatever the function is.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

explain please

OpenStudy (solomonzelman):

yes, so for example if we have: \(\large\color{black}{ R(x)=x^3-2x^2-x+2}\) then you set: \(\large\color{black}{ 0=x^3-2x^2-x+2}\) ,and solve for x.

OpenStudy (solomonzelman):

\(\large\color{black}{ 0=x^3-2x^2-x+2}\) \(\large\color{black}{ 0=x^2(x-2)-(x-2)}\) \(\large\color{black}{ 0=(x^2-1)(x-2)}\) \(\large\color{black}{ 0=(x-1)(x+1)(x-2)}\) so your x-intercepts would be (in my example) \(\large\color{black}{ 1,~-1,~~~2}\) .

OpenStudy (solomonzelman):

so you would mark the points \(\large\color{black}{ (-1,0)}\) , \(\large\color{black}{ (1,0)}\) , \(\large\color{black}{ (2,0)}\) .

OpenStudy (solomonzelman):

in your example I could usually imagine a very high degree function. so you would have many x-intercepts, but this is ain't important

OpenStudy (solomonzelman):

you are: ~ Setting R(x)=0 (replacing R(x) by a 0 ) ~ Solve for x ~ Write the solutions as points)

OpenStudy (anonymous):

ok let me try to understand

OpenStudy (solomonzelman):

sure, take your time

OpenStudy (anonymous):

R(x)=x3-2x2x+2 set R(x) to 0 0=x3-2x2-x+2 0=x2(x-4)(x-4) 0=(x2-2)(x-2) 0=(x-2)(x+2)(x-4)

OpenStudy (anonymous):

@SolomonZelman is this correct so far?

OpenStudy (solomonzelman):

were you doing, \(\large\color{black}{ R(x)=x^3-2x^2-x+2 }\) ?

OpenStudy (solomonzelman):

you have a little error I think

OpenStudy (anonymous):

yes i am and ok help me please

OpenStudy (solomonzelman):

\(\large\color{black}{ R(x)=x^3-2x^2-x+2 }\) \(\large\color{black}{ 0=x^3-2x^2-x+2 }\) good till now, then you factor: \(\large\color{black}{ R(x)=x^3-2x^2-x+2 }\) \(\LARGE\color{white}{ \rm \left| \right| }\) \(\large\color{black}{ 0=x^2(x-2)-(x-2) }\) \(\LARGE\color{white}{ \rm \left| \right| }\) \(\large\color{black}{ 0=x^2(x-2)-1(x-2) }\) \(\LARGE\color{white}{ \rm \left| \right| }\) \(\large\color{black}{ 0=(x^2-1)(x-2) }\) \(\LARGE\color{white}{ \rm \left| \right| }\) \(\large\color{black}{ 0=(x^2-1^2)(x-2) }\) \(\LARGE\color{white}{ \rm \left| \right| }\) \(\large\color{black}{ 0=(x+1)(x-1)(x-2) }\) the the zeros are: -1, 1, 2 \(\LARGE\color{white}{ \rm \left| \right| }\)

OpenStudy (solomonzelman):

you don't have to use my function (that I offered as an example) though...

OpenStudy (anonymous):

ok thanks i will post another one

OpenStudy (solomonzelman):

sure.... (but remember, one question per thread) yw, and good luck!

OpenStudy (anonymous):

i know

OpenStudy (solomonzelman):

god, yw 1s again

OpenStudy (adrianna.gongora):

What was the answer ? ;-;